OpenStudy (anonymous):

Mario Agnello is opening a new pizza shop in town. He plans to offer a twelve -inch diameter pizza and an eighteen -inch diameter pizza. He has set the price of a twelve -inch cheese pizza at eight dollars , based on the amount of ingredients needed to cover that size crust. If Mario would like to keep the pricing proportional, what should he charge for an eighteen -inch cheese pizza? Show and explain your solution.

OpenStudy (mathmagician):

Area(18)=3.14*18^2=1017.36 Area(12)=3.14*12^2=452.16 452.16=8 dollars 1017.36-x x=8*1017.36/452.16=18 dollars

OpenStudy (anonymous):

:O

OpenStudy (anonymous):

If I use that answer, my teacher will probaly know i am cheating cause i am not that smart